For the given probability distribution function

X=x 1. 2 3. 4. 5

P[X=x] 0.1 0.3 0.2 0.2 0.2


E(x) =?

E(2x+5) =?

E(x/2+3)=?

E(x 2 )


1
Expert's answer
2021-01-12T16:20:20-0500

Find E(X):


solution :


we need to multiply the corresponding X outcomes with the corresponding probabilities, in order to compute the population mean μ.




Therefore, the population mean is calculated as follows:

"\u03bc\n\u200b\t\n \n=\n\n\u200b\t\n \n\u2211 \n\u200b\t\n X \n\n *p(X \ni\n\u200b\t\n )"

=1⋅0.1+2⋅0.3+3⋅0.2+4⋅0.2+5⋅0.2

=3.1

Hence E(X)= 3.1


Find E(2x+5):

Solution :

We know E[aX + b] = aE[X] + b

Hence E(2X+5)= 2E(X)+5

=2(3.1)+5

=11.2

find E(X/2+3)= "\frac{1}{2}E(X)+3"

="\frac{1}{2}(3.1)+3"

=4.55

find "E(X^2):"

solution:

we need to multiply the corresponding squares "X^2" of the outcomes with the corresponding probabilities, in order to compute "E(X^2)" :




"E(X ^\n2\n )\n\u200b\t\n \n=\n \n\u2211 \n\t\n X \n^\n2\n\u200b*\t\n p(X \n\n\u200b)"


="1 ^\n2\n \u22c50.1+2 ^\n2\n \u22c50.3+3 ^\n2\n \u22c50.2+4 ^\n2\n \u22c50.2+5 ^\n2\n \u22c50.2=\n11.3"